OpenStudy (anonymous):

find the area of enclosed by the gievn curve. r= a cos 2(θ)

OpenStudy (anonymous):

Look at this similar example have just solved: Find the area enclosed by the curve: r=2+3cosθ Solution: Area of curve, A=2∫arccos(−2/3)0r22dϕ ⟹A=2∫arccos(−2/3)0(2+3cosϕ)22dϕ =2∫arccos(−2/3)04+9cos2ϕ+12cosϕ2dϕ =∫arccos(−2/3)0(4+9cos2ϕ+12cosϕ)dϕ =35√+172cos−1(−23). Does it make some sense to you?

OpenStudy (anonymous):

There is a symmetry about x-axis.Do u think so?

OpenStudy (anonymous):

If \(r=a\cos^2\theta\), then the area would be \[A=\frac{1}{2}\int_a^br^2~d\theta=\frac{1}{2}\int_0^{2\pi}\left(2\cos^2\theta\right)^2~d\theta\\ A=2\int_0^{2\pi}\cos^4\theta~d\theta\] Use the half-angle identity: \[\cos^{2}x=\frac{1}{2}\left(1+\cos2x\right)\] So you have \[\cos^{4}x=\left(\cos^2x\right)^2=\left(\frac{1}{2}\left(1+\cos2x\right)\right)^2\]